Mathematics 242 – Principles of Analysis Definitions Quiz 2 – September 16, 2004

Your Name:

1. What does it mean to say a real number s is the supremum (or least upper bound) of a set A of real numbers: s = sup(A).

2. What is sup ({x ∈ Q : x^{2} < 3})?